# MunchSonic: Tracking Fine-grained Dietary Actions through Active Acoustic Sensing on Eyeglasses

### v1

@@ Abstract.
We introduce MunchSonic, an AI-powered active acoustic sensing system integrated into eyeglasses, designed to track fine-grained dietary actions such as hand-to-mouth movements for food intake, chewing, and drinking. MunchSonic emits inaudible ultrasonic waves from a commodity eyeglass frame. The reflected signals contain detailed information about the positions and movements of various body parts, including the mouth, jaw, arms, and hands, all involved in eating activities. These signals are processed by a custom deep learning pipeline to classify six actions: food intake, chewing, drinking, talking, face-hand touching, and other activities not belonging to the tracking set (null). In an unconstrained user study with 12 participants, MunchSonic achieved a 93.5% macro F1-score in a user-independent evaluation with a 2-second time resolution, demonstrating its effectiveness. Additionally, MunchSonic accurately tracks eating episodes and the frequency of food intake within those episodes.
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 2024 International Symposium on Wearable Computers; October 8–12, 2024; Melbourne, Australiabooktitle: Proceedings of the 2024 International Symposium on Wearable Computers (ISWC ’24), October 5–9, 2024, Melbourne, Australiaprice: 15.00doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/XX/XX
[…]
Research
@@ 1. Introduction
To address this challenge, we developed MunchSonic, a fine-grained dietary action recognition system in the form factor of commodity eyeglasses using active acoustic sensing. Similar to ActSonic (Mahmud et al., 2024), our system employs two pairs of microphones and speakers attached to the hinges of the eyeglasses. MunchSonic transmits inaudible ultrasonic waves that reflect off various body parts, generating a two-dimensional echo profile, which serves as a range spectrogram of the scanned body parts. This profile is then analyzed by a lightweight deep learning framework to infer fine-grained dietary actions. Our system is based on the observation that fine-grained dietary actions involve movements of the mouth, jaw, and arms, which can be inferred from reflected ultrasound, as shown in prior work (Mahmud et al., 2023; Li et al., 2022, 2024). Therefore, we believe it is possible to distinguish these actions from the reflected signals beyond merely identifying eating or drinking moments (Mahmud et al., 2024). We evaluated our system in a user study with 12 participants in unconstrained environments of their choice. The results showed that MunchSonic can accurately recognize fine-grained dietary actions such as hand movements for intake, chewing, and drinking, as well as non-dietary actions like talking, face touching, and other activities (null), achieving an F1-score of 93.5%.
@@ 2. Related Work
2015 Thomaz et al. (Thomaz et al., 2015b)
Wrist-mounted
Device
Microphone
(Passive Audio)
Meals F1-score = 79.8% -
2016 Farooq et al. (Farooq and Sazonov, 2016a) Eyeglasses Piezoelectric
Food
Intake
F1-score = 100.0%* -
2016 iHear Food (Gao et al., 2016)
Bluetooth
Headset
Microphone
(Passive Audio)
Eating
Episodes
Accuracy = 76.82% -
@@ 1 Proximity
Table 1. Comparison of eating detection studies based on wearable form factors, sensing modalities, granularity of detected events, evaluation metrics (F1-score or precision/recall), and power draw (in mW). Asterisks (*) in the Evaluation Metric column indicate that the system was evaluated in lab studies; otherwise, the evaluation was conducted in free-living conditions.
@@
Various wearable devices with different sensing modalities have been proposed to track eating events. This section first discusses non-eyeglass wearables, followed by eyeglasses. Table 1 summarizes prior studies on tracking eating activities.
@@
Form Factors Other than Eyeglasses. Sensing systems for detecting eating behaviors have been integrated into various wearable devices. Smartwatch-based systems for eating (Thomaz et al., 2015a; Ye et al., 2016; Dong et al., 2013; Kyritsis et al., 2019; Sen et al., 2018; Zhang et al., 2022) and drinking (Hamatani et al., 2018), worn on the user’s dominant hand, mainly use Inertial Measurement Units (IMUs) to track hand-to-mouth gestures. These systems face challenges in distinguishing similar hand motions, leading to a high incidence of false positives in real-world applications (Tong et al., 2020; Chun et al., 2018a). Ear-worn devices, or earables, are also used to track eating events due to their proximity to the jaw, which moves during chewing. Techniques such as in-ear proximity sensing (Bedri et al., 2015a, b), passive acoustic sensing (Gao et al., 2016; Bi et al., 2018), or a combination of both (Bedri et al., 2017), have been employed to monitor chewing activity. Neckbands (Farooq et al., 2014; Rahman et al., 2014; Yatani and Truong, 2012) and necklaces (Alshurafa et al., 2015; Kalantarian et al., 2015; Chun et al., 2018b; Shin et al., 2019; Zhang et al., 2020b) use passive acoustic and proximity sensing to track sounds and motions associated with chewing and swallowing. Despite these advancements, necklaces, and neckbands have not gained the same popularity as other wearables (Amores et al., 2019).
@@
Eyeglasses Form Factor. Dietary monitoring using eyeglasses is of particular interest because sensors located near the mouth and jaw can capture significant movement during eating and drinking. Additionally, eyeglasses have higher social acceptance (Kosmyna et al., 2019) (64% adoption among the US population (Farooq and Sazonov, 2018)) compared to other wearables. Researchers have attached Electromyography (EMG) electrodes to eyeglasses to monitor eating behaviors (Zhang and Amft, 2017). Piezoelectric sensors (Farooq and Sazonov, 2016a, b), which make skin contact, have been used to track jaw movements but the performance tends to be impacted under conditions of sweating or rapid user movement. Inertial sensors (Farooq and Sazonov, 2018) and load cells (Chung et al., 2017) have been proposed for dietary event tracking but have not shown promising results. Sensor fusion approaches (Bedri et al., 2020; Rahman et al., 2016; Mirtchouk et al., 2017; Bedri et al., 2022; Shin et al., 2022), integrating gyroscopes, accelerometers, proximity sensors, and microphones on eyeglasses, have shown performance improvements in tracking eating events. Passive acoustic sensing using contact microphones (Bi et al., 2018) captures chewing sounds but requires close contact between the sensor and the skin. Acoustic sensing-based activity recognition systems have included labels for eating (Mahmud et al., 2024) or drinking (Mollyn et al., 2022; Mahmud et al., 2024), but they focus on eating episodes rather than fine-grained actions.
@@
In summary, compared to these prior work, MunchSonic is the first to infer fine-grained dietary actions by employing active acoustic sensing on eyeglasses to capture information related to fine-grained jaw and hand movements. This approach allows for detailed dietary action extraction, such as intake and chewing, unlike other systems summarized in Table 1, which primarily track the moments of eating activities.
@@ 3. System Implementation
The goal of MunchSonic is to capture the information related to movements on multiple body parts involved in dietary activities using an eyeglasses form factor. Previous research has shown active acoustic sensing’s efficacy in tracking facial muscle movements (Li et al., 2022, 2024; Zhang et al., 2023), upper body limbs (Mahmud et al., 2023), respiration (Wang et al., 2018), sign language gestures (Jin et al., 2021, 2023), and everyday activities (Mahmud et al., 2024). Inspired by these prior work, especially (Mahmud et al., 2024), MunchSonic integrates active acoustic sensing into eyeglasses to track these fine-grained dietary actions. This section discusses the active acoustic sensing mechanism, the hardware implementation, and the deep learning framework for processing the captured data.
@@ 3.1. Sensing Mechanism
The design of MunchSonic’s sensing system is inspired by the active acoustic sensing approach from (Mahmud et al., 2024). In essence, MunchSonic uses a similar active acoustic sensing processing method, based on Cross-correlation based Frequency Modulated Continuous Wave (C-FMCW) (Wang et al., 2018). These chirps transmitted from eyeglasses’ hinges, with frequency ranges of 18-21.5 KHz and 21.5-24.5 KHz for the left and right transmitters, respectively. The receiver samples at 50 KHz, with each chirp containing 600 samples and a sweep period of 0.012 seconds. The system can detect changes as small as 3.43 mm, given the slow speed of mouth, jaw, and hand movements compared to the speed of sound (343 m/s). The sensing range is up to 2.058 meters, though a shorter range is used for fine-grained dietary actions.
@@ 3.2. Hardware and Form Factor
We replicated the sensing system from ActSonic (Mahmud et al., 2024) with minor changes to the sensors and controller unit, using the OWR-05049T-38D speaker and ICS-43434 microphone (TDK, [n. d.]). Our customized controller unit, shown in Figure 1(b), features the nRF52840 microcontroller (Semiconductors, [n. d.]), two MAX98357A audio amplifiers, a BLE SGW1110 (SGWireless, [n. d.]) module, power management modules, and an SD card slot with a SanDisk Extreme Pro microSD card (Amazon.com, [n. d.]), optimized for power efficiency. The transceiver boards, each with a speaker and microphone, are attached to the eyeglass hinges, with placement refined for optimal tracking of mouth, jaw, and hand movements. The controller unit and LiPo battery are attached to one leg of the glasses frame, connected via Flexible Printed Circuit (FPC) cables and a JST connector. The system draws 96.5 mW of power, with a voltage of 4.02V and current of 24.0 mA, and the prototype used in the MunchSonic study, featuring a 290 mAh LiPo battery, lasts approximately 11.25 hours in SD card storage mode.
@@ 3.3.1. Data Processing and Model Architecture
We designed a lightweight deep learning framework to process the active acoustic data captured by the MunchSonic system. The differential echo profile described in 3.1 serves as the input to the deep learning model. We create overlapping sliding windows from the stream of differential echo profile data. To augment the data, we apply Gaussian noise to 5% of the sliding windows and mask 5% of the distance axis ($y$-axis) of the sliding window to simulate environmental noise and unrelated movements.
The preprocessed sliding windows are fed into a MobileNetV2 (Sandler et al., 2018) convolutional neural network encoder. This lightweight encoder, effective for edge devices, generates a 256-dimensional embedding vector for each sliding window. This vector is then fed into a classifier network with three feedforward layers (128, 64, and 6 neurons respectively). Each layer, except the last one, includes batch normalization (Ioffe and Szegedy, 2015), dropout (Srivastava et al., 2014) with a probability of 0.25, and Leaky ReLU (Xu et al., 2015) activation. Finally, a softmax operation provides the class probability distribution.
@@ 3.3.2. Model Training
The input to MunchSonic’s deep learning framework consists of sliding windows of differential echo profiles. The optimal sliding window for the MunchSonic dataset is 2 seconds long with 50% overlap. This window length was chosen because it appropriately fits an eating intake gesture. The optimal sensing range is 150 pixels on the $y$-axis of the differential echo profile, corresponding to 51.45 cm from the MunchSonic device. With speakers and microphones co-located on each hinge of the eyeglasses and transmitting acoustic waves in two frequency ranges, there are four channels in the differential echo profile (two direct and two indirect paths). Thus, the shape of the input sliding window is $(4\times 150\times 166)$, as each second of the sliding window contains $\lfloor\frac{50000}{600}\rfloor=83$ samples.
We use focal loss (Lin et al., 2017) for training the MunchSonic model due to its effectiveness in handling class imbalance in the MunchSonic dataset. The Adam (Kingma and Ba, 2014) optimizer with a cosine annealing learning rate scheduler is employed, starting with an initial learning rate of $10^{-2}$. The model, implemented using PyTorch and PyTorch Lightning frameworks, is trained for 30 epochs with a batch size of 128 on GeForce RTX 2080 Ti GPUs.
@@ 4. User Study
We conducted a 45-minute user study to evaluate MunchSonic in an unconstrained setting. This section details the study design and the resulting dataset.
4.1. Design of User Study
The active acoustic sensing-based activity recognition system ActSonic (Mahmud et al., 2024) has proven effective in tracking eating and drinking episodes in real-world settings over long-duration studies. For example, it can identify eating and drinking moments with a one-second resolution, achieving F1 scores of 90% and 88% respectively, without needing training data from new users in new environments. Building on this capability, the MunchSonic user study was designed with the assumption that active acoustic sensing on glasses can already accurately distinguish eating moments from other activities. This paper focuses on further extracting fine-grained dietary actions, such as hand-to-mouth movements for food intake, chewing, and drinking, as well as distinguishing eating from non-eating actions from active acoustic signals once eating moments are recognized. Specifically, the MunchSonic user study aimed to evaluate the system’s performance in unconstrained conditions in tracking fine-grained dietary actions and differentiating these from non-eating actions. We adopted a relatively shorter-duration free-living study, leveraging ActSonic’s proven efficacy in longer-duration in-the-wild studies. Additionally, ground truth annotation in our study was done per second, unlike other eating detection systems (Shin et al., 2022) that rely on self-reporting, enabling us to recognize dietary actions with a high time resolution of two seconds.
Figure 2. Foods consumed by user study participants.
4.2. Study Protocol
We conducted a user study approved by our organization’s Institutional Review Board (IRB) with 13 participants, 6 identifying as male and 7 as female, with an average age of $26.5$ years, ranging from 21 to 36 years. The participants wore the MunchSonic eyeglasses form factor and a chest-mounted GoPro HERO9 camera (GoPro, 2020) facing upward to capture their eating activities around the face as the ground truth. The camera recorded 720p video at 30 fps with a diagonal field of view of $148^{\circ}$. Note that one participant (P09) accidentally turned off the chest-mounted camera, resulting in no ground truth data for that participant. Therefore, we discarded that participant from the study and evaluated the system on the remaining 12 participants.
@@ 4.2. Study Protocol
4.3. Dataset Statistics
We collected 540 minutes of data from 12 participants. Two researchers labeled the reference video from the chest-mounted camera at each second using the ANU-CVML Video Annotation Tool (Vidat) (Zhang et al., 2020a). The data was categorized into six classes: hand-to-mouth movement for food intake, chewing, drinking, and talking, face touch, and a null class for other activities. A synchronization script was used to extrapolate ground truth labels for the differential echo profile sliding windows. The class distribution was 53.5% null, 5.9% hand-to-mouth for food intake, 22.0% chewing, 2.2% drinking, 15.1% talking, and 1.3% touching face with hand for non-eating activities.

### v2

@@ Abstract.
We introduce MunchSonic, an AI-powered active acoustic sensing system integrated into eyeglasses to track fine-grained dietary actions. MunchSonic emits inaudible ultrasonic waves from the eyeglass frame, with the reflected signals capturing detailed positions and movements of body parts, including the mouth, jaw, arms, and hands involved in eating. These signals are processed by a deep learning pipeline to classify six actions: hand-to-mouth movements for food intake, chewing, drinking, talking, face-hand touching, and other activities (null). In an unconstrained study with 12 participants, MunchSonic achieved a 93.5% macro F1-score in a user-independent evaluation with a 2-second resolution in tracking these actions, also demonstrating its effectiveness in tracking eating episodes and food intake frequency within those episodes.
Eating Detection; Acoustic Sensing; Activity Recognition
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 2024 ACM International Symposium on Wearable Computers; October 5–9, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 2024 ACM International Symposium on Wearable Computers (ISWC ’24), October 5–9, 2024, Melbourne, VIC, Australiadoi: 10.1145/3675095.3676619isbn: 979-8-4007-1059-9/24/10ccs: Human-centered computing Ubiquitous and mobile computing systems and tools
@@ 1. Introduction
To address this challenge, we developed MunchSonic, a fine-grained dietary action recognition system in the form factor of commodity eyeglasses using active acoustic sensing. Similar to ActSonic (Mahmud et al., 2024), our system employs two pairs of microphones and speakers attached to the hinges of the eyeglasses. MunchSonic transmits inaudible ultrasonic waves that reflect off various body parts, generating a two-dimensional echo profile, which serves as a range spectrogram of the scanned body parts. This profile is then analyzed by a lightweight deep learning framework to infer fine-grained dietary actions.
MunchSonic is based on the observation that fine-grained dietary actions involve movements of the mouth, jaw, and arms, which can be inferred from reflected ultrasound, as shown in prior work (Mahmud et al., 2023; Li et al., 2022; eyeEcho; Jin et al., 2021). However, the capacity of the active acoustic sensing-based general-purpose activity recognition system ActSonic (Mahmud et al., 2024) is limited to tracking moments of eating and drinking episodes. Since eating is a complex activity involving several fine-grained dietary actions such as hand-to-mouth movement for intake, chewing, swallowing, etc., we designed the MunchSonic system specifically for tracking body parts involved in eating, with modifications to ActSonic (Mahmud et al., 2024) hardware and inference pipeline. We evaluated our system in a user study with 12 participants in unconstrained environments of their choice. The results showed that MunchSonic can accurately recognize fine-grained dietary actions such as hand movements for intake, chewing, and drinking, as well as non-dietary actions like talking, face touching, and other activities (null), achieving an F1-score of 93.5%.
@@ 1 Proximity
Table 1. Comparison of eating detection studies based on wearable form factors, sensing modalities, event granularity, evaluation metrics (F1-score or precision/recall), and power draw (mW). Asterisks (*) in the Evaluation Metric column indicate lab study evaluations; otherwise, evaluations were in free-living conditions. The dagger symbol (†) in the Power Consumption column indicates machine learning inference was performed on a cloud or mobile device, not the wearable.
@@
Various wearable devices with different sensing modalities have been proposed to track eating events. This section discusses non-eyeglass wearables, followed by eyeglasses. Table 1 summarizes prior studies on tracking eating activities.
@@
Form Factors Other than Eyeglasses. Smartwatch-based systems for eating (Thomaz et al., 2015a; Ye et al., 2016; Dong et al., 2013; Kyritsis et al., 2019; Sen et al., 2018; Zhang et al., 2022; Thomaz et al., 2015b) and drinking (Hamatani et al., 2018) use IMUs to track hand-to-mouth gestures but face challenges distinguishing similar motions, leading to high false positives (Tong et al., 2020; Chun et al., 2018a; Mahmud et al., 2020; Tonmoy et al., 2021). Ear-worn devices, or earables, track eating by monitoring jaw movement through in-ear proximity sensing (Bedri et al., 2015a, b), passive acoustic sensing (Gao et al., 2016; Bi et al., 2018), or both (Bedri et al., 2017). Neckbands (Farooq et al., 2014; Rahman et al., 2014; Yatani and Truong, 2012) and necklaces (Alshurafa et al., 2015; Kalantarian et al., 2015; Chun et al., 2018b; Shin et al., 2019; Zhang et al., 2020b) use passive acoustic and proximity sensing to track chewing and swallowing. Despite advancements and around $80\%$ accuracy in free-living conditions, neckband and necklace form factors have not gained the same popularity as other wearables (Amores et al., 2019).
@@
Eyeglasses Form Factor. Dietary monitoring using eyeglasses is promising due to their proximity to the mouth and jaw and higher social acceptance (Kosmyna et al., 2019) (64% adoption in the US (Farooq and Sazonov, 2018)). Researchers have used EMG electrodes (Zhang and Amft, 2017), piezoelectric sensors (Farooq and Sazonov, 2016a, b), inertial sensors (Farooq and Sazonov, 2018), and load cells (Chung et al., 2017) on eyeglasses to track eating behaviors. Sensor fusion approaches (Bedri et al., 2020; Rahman et al., 2016; Mirtchouk et al., 2017; Bedri et al., 2022; Shin et al., 2022), integrating gyroscopes, accelerometers, proximity sensors, and microphones, have shown improved performance. Passive acoustic sensing using contact microphones (Bi et al., 2018) captures chewing sounds but requires close skin contact. Systems using acoustic sensing (Mahmud et al., 2024; Mollyn et al., 2022) focus on eating episodes rather than fine-grained actions.
@@
In summary, MunchSonic is the first to use active acoustic sensing on eyeglasses to infer fine-grained dietary actions by capturing jaw and hand movements. This allows for detailed dietary action extraction, such as intake and chewing, unlike other systems that primarily track moments of eating episodes.
@@ 3. System Implementation
The goal of MunchSonic is to capture the information related to movements on multiple body parts involved in dietary activities using an eyeglasses form factor. Previous research has shown active acoustic sensing’s efficacy on wearables in tracking facial muscle movements (Li et al., 2022, 2024a; Zhang et al., 2023b, a; Sun et al., 2023), upper body limbs (Mahmud et al., 2023), hand poses (Lee et al., 2024; Yu et al., 2024), gaze (Li et al., 2024b), respiration (Wang et al., 2018), sign language gestures (Jin et al., 2021, 2023), physiological signals (Fan et al., 2023) and everyday activities (Mahmud et al., 2024). Inspired by these prior works, especially (Mahmud et al., 2024), MunchSonic integrates active acoustic sensing into eyeglasses to track these fine-grained dietary actions. This section discusses the active acoustic sensing mechanism, the hardware implementation, and the deep learning framework for processing the captured data.
@@ 3.1. Sensing Mechanism
The design of MunchSonic’s sensing system is inspired by the active acoustic sensing approach from (Mahmud et al., 2024). In essence, MunchSonic uses a similar active acoustic sensing processing method, based on cross-correlation-based Frequency Modulated Continuous Wave (C-FMCW) (Wang et al., 2018). These chirps are transmitted from eyeglasses’ hinges, with frequency ranges of 18-21.5 KHz and 21.5-24.5 KHz for the left and right transmitters, respectively. The receiver samples at 50 KHz, with each chirp containing 600 samples and a sweep period of 0.012 seconds. The system can detect changes as small as 3.43 mm, given the slow speed of mouth, jaw, and hand movements compared to the speed of sound (343 m/s). The sensing range is up to 2.058 meters, though a shorter range is used for fine-grained dietary actions. MunchSonic uses a lower amplifier gain compared to (Mahmud et al., 2024) to focus on mouth, jaw, and hand movements near the face for tracking fine-grained dietary actions. Consequently, the microphones receive weaker reflections from lower body parts that are not involved in dietary actions.
@@ 3.2. Hardware and Form Factor
We replicated the sensing system from ActSonic (Mahmud et al., 2024) with minor changes to the sensors and controller unit, using the OWR-05049T-38D speaker and ICS-43434 microphone (TDK, [n. d.]). Our customized controller unit, shown in Figure 1(b), features the nRF52840 microcontroller (Semiconductors, [n. d.]), two MAX98357A audio amplifiers, a BLE SGW1110 (SGWireless, [n. d.]) module, power management modules, and an SD card slot with a SanDisk Extreme Pro microSD card (Amazon.com, [n. d.]), optimized for power efficiency. The transceiver boards, each with a speaker and microphone, are attached to the eyeglass hinges, with placement refined for optimal tracking of mouth, jaw, and hand movements. The controller unit and LiPo battery are attached to one leg of the glasses frame, connected via Flexible Printed Circuit (FPC) cables and a JST connector. The sensing system on the eyeglasses draws 96.5 mW of power for collecting data, with a voltage of 4.02V and a current of 24.0 mA. The prototype used in the MunchSonic study, featuring a 290 mAh LiPo battery, lasts approximately 11.25 hours in SD card storage mode.
@@ 3.3.1. Data Processing and Model Architecture
We designed a lightweight deep learning framework to process the active acoustic data captured by the MunchSonic system. The differential echo profile described in Section 3.1 serves as the input to the model. We create overlapping sliding windows from the differential echo profile data. To augment the data, we apply Gaussian noise to 5% of the windows and mask 5% of the distance axis to simulate environmental noise and unrelated movements.
The preprocessed sliding windows are fed into a MobileNetV2 (Sandler et al., 2018) convolutional neural network encoder, which generates a 256-dimensional embedding vector for each window. This vector is then fed into a classifier network with three feedforward layers (128, 64, and 6 neurons). Each layer, except the last, includes batch normalization (Ioffe and Szegedy, 2015), dropout (Srivastava et al., 2014) with a probability of 0.25, and Leaky ReLU (Xu et al., 2015) activation. Finally, a softmax operation provides the class probability distribution.
@@ 3.3.2. Model Training
The input to MunchSonic’s deep learning framework consists of sliding windows of differential echo profiles, optimally 2 seconds long with 50% overlap. This window length fits an eating intake gesture. The optimal sensing range is 150 pixels on the $y$-axis of the differential echo profile, corresponding to 51.45 cm from the MunchSonic device. With speakers and microphones on each hinge of the eyeglasses, transmitting acoustic waves in two frequency ranges, there are four channels in the differential echo profile. Thus, the shape of the input window is $(4\times 150\times 166)$, as each second contains 83 samples.
We use focal loss (Lin et al., 2017) for training to handle class imbalance in the MunchSonic dataset. The Adam (Kingma and Ba, 2014) optimizer with a cosine annealing learning rate scheduler starts with an initial learning rate of $10^{-2}$. The model, implemented using PyTorch and PyTorch Lightning, is trained for 30 epochs with a batch size of 128 on GeForce RTX 2080 Ti GPUs.
@@ 4. User Study
We conducted a 45-minute user study to evaluate MunchSonic in an unconstrained setting. Building on ActSonic’s (Mahmud et al., 2024) success in tracking eating and drinking episodes with high F1 scores, the MunchSonic study assumed active acoustic sensing on glasses can distinguish eating moments from other activities. This paper focuses on extracting fine-grained dietary actions, such as hand-to-mouth movements, chewing, and drinking. The study aimed to assess the system’s performance in tracking these actions and distinguishing them from non-eating activities in real-world conditions. We conducted a shorter-duration free-living study, leveraging ActSonic’s proven efficacy in longer studies. Ground truth annotation was done per second, unlike other systems (Shin et al., 2022) that rely on self-reporting, enabling high-resolution recognition of dietary actions.
Study Protocol We conducted a user study approved by our organization’s Institutional Review Board (IRB) with 13 participants, 6 identifying as male and 7 as female, with an average age of $26.5$ years, ranging from 21 to 36 years. The participants wore the MunchSonic eyeglasses form factor and a chest-mounted GoPro HERO9 camera (GoPro, 2020) facing upward to capture their eating activities around the face as the ground truth. The camera recorded 720p video at 30 fps with a diagonal field of view of $148^{\circ}$. Note that one participant (P09) accidentally turned off the chest-mounted camera, resulting in no ground truth data for that participant. Therefore, we discarded that participant from the study and evaluated the system on the remaining 12 participants.
Figure 2. User study in unconstrained conditions: (a) Foods consumed by user study participants, (b) Sample images of the chest-mounted camera view of MunchSonic data collection pipeline.
@@
Dataset Statistics We collected 540 minutes of data from 12 participants. Two researchers labeled the reference video from the chest-mounted camera at each second using the ANU-CVML Video Annotation Tool (Vidat) (Zhang et al., 2020a). The inter-rater reliability was measured through Cohen’s Kappa (Lazar et al., 2017), $\kappa$, and its value is 0.82 ($\kappa>0.80$ denotes near perfect agreement, and $0.6<\kappa\leq 0.8$ denotes satisfactory agreement). The camera view that the annotators labeled is shown in Figure 2(b). The data was categorized into six classes: hand-to-mouth movement for food or drink intake, chewing, drinking, talking, face touch, and a null class for other activities. The annotators were instructed to segment the actions in two phases. First, they determined whether the action was eating or non-eating. If the action was identified as eating, they further categorized it into fine-grained actions such as hand-to-mouth movement, chewing, and drinking. Swallowing was integrated into the chewing category. For non-eating actions, they segmented the activities into talking, face touching, and other activities not included in the tracking set. A synchronization script was used to extrapolate ground truth labels for the differential echo profile sliding windows. The class distribution was 53.5% null, 5.9% hand-to-mouth for food intake, 22.0% chewing, 2.2% drinking, 15.1% talking, and 1.3% touching face with hand for non-eating activities.